Geometry: Find the dimensions of the rectangle shown at the right if the length is twice it's width.
________________
| |
| | The Perimeter is 208
----------------

Answer by TutorDelphia(193)

We are told the length is twice the width which we write as:
l=2w
We also know the perimeter (which is all the sides added, the length plus the width plus the length plus the width) is 208 which we can write as:
2l+2w=208
Now we just sub in for the length with 2w
2(2w)+2w=208
distribute
4w+2w=208
combine like terms
6w=208
divide by 6
w=34 2/3
sub 34 2/3 back in to l=2w
l=69 1/3